A single die is rolled twice. Find the probability of rolling an odd number the first time and a number greater than 3 the second time.
Having trouble with the problem.

Answer by Fombitz(32388)

P(odd)=3/6=1/2 (1,3,5)
P(>3)=3/6=1/2 (4,5,6)
The events are independent, multiply the probabilities.
P(odd then >3)=(1/2)(1/2)=1/4